In a certain town 60% of the households own mutual funds, 40% own individual stocks, and 20% own both mutual funds and individual stocks.

1.         The proportion of households that own mutual funds but not individual stocks is

            a.   20%.

            b.   30%.

            c.   40%.

            d.   50%.

2.         The proportion of households that own neither mutual funds nor individual stocks is

            a.   20%.

            b.   30%.

            c.   40%.

            d.   50%.

3.         The proportion of households that own mutual funds but not individual stocks or individual stocks but not mutual funds is

            a.   40%.

            b.   60%.

            c.   80%.

            d.   100%.
Answer #1

P(M)= 0.6 PLI) 2 0.4 pcmo I) - 0.2 Pima- pimos) = 0.6-0.2 = 0.4 = 40V. in Plown only mutual funds) 2) p (neither mutal funds